# MCHex: Marching Cubes Based Adaptive Hexahedral Mesh Generation with Guaranteed Positive Jacobian

### v2

@@ 4 Properties of 3-regular graph
This section begins by establishing the definitions of a $k$-regular graph, a $k$-regular polyhedron, and the midpoint subdivision operation. It then provides a formal proof that the midpoint subdivision of a polyhedron induces all-hex elements if and only if the polyhedron is 3-regular. The necessity of this condition is also proved by showing that any 3-regular polyhedron satisfy the Global Necessary Condition for hex meshing proposed in (liu2018singularity).
@@
As an extension, it holds that the three-dimensional CW complex $P^{\prime}=(V^{\prime},E^{\prime},F^{\prime},C^{\prime})$ induced by 3-regular two-dimensional CW complex $P=(V,E,F)$ satisfy the hex-meshing Global Necessary Condition (liu2018singularity). This equation establishes a global constraint that the resulting hex mesh in $P^{\prime}$ must satisfy:
@@ 5 Positive Jacobian Proof
This section begins by defining the Jacobian determinant as a key measure of element quality. The MC method (chernyaev1995marching) is then applied to reconstruct the isosurface within each cell $h\in H_{\text{Int}(T_{\text{AABB}})}$ where the SDF values at its eight grids exhibit sign changes. It is demonstrated that the cells intersected by the isosurface are all 3-regular, making them hex meshable via midpoint subdivision. Among a total of thirty-one topological cases in a marching cube, fourteen templates guaranteed to generate positive Jacobian hexes under midpoint subdivision are listed and proven. The other seventeen cases for which a guarantee failed to be established are transferred into the fourteen templates using local refinement. Local refinement is introduced to transform these failed cases into the fourteen cases with proof certificates. Additionally, since the cell intersected by the isosurface can be transition cells or non-cubic cells, the positive Jacobian proofs are therefore given to not only the cubic cell, but all possible deformed hex elements. As a proof of concept, complete proofs for the vanilla 3-refinement templates presented in (schneiders2000octree, zhang2006adaptive) are provided. Based on the experience, it is highly feasible to extend the proofs to element-saving 3-refinement templates (ito2009octree) and 2-refinement templates (zhang2013robust, owen2017template) as well. Such tasks are left for future work.
@@
While a bijective mapping requires a positive Jacobian determinant $\lvert\mathbf{J}\rvert>0$ everywhere within an element, it remains an open problem to check the global minimum in the parameter space. An alternative approach, popularized by the visualization tool ParaView (ayachit2015paraview) and recent literature (bracci2019hexalab, gao2019feature, guo2020cut), is to evaluate the Jacobian at nine points: the eight corners and the element center in the parametric space. The minimum value from these nine points is then used as a metric for element quality. The objective of the following analysis is to provide guarantees for positivity at these nine points.
@@
This proof extends the univariate case established in (mathar2018orthogonal) Appendix B to the multivariate setting. Theorem 5.1 tells that the coefficients $\beta_{\mathbf{i}}$ of the Bernstein basis are related to the monomial coefficients $J_{\mathbf{i}}$ via the linear transformation.
@@
Given an all-hex tessellation of the bounding box $T_{\text{AABB}}$ and the SDF value at eight corners of each cell $h\in H_{\text{Int}(T_{\text{AABB}})}$ with respect to the input geometry $T_{\text{geom}}$, the MC algorithm reconstructs an isosurface. An edge is a sign-change edge if its two endpoints lie on opposite sides of the isosurface (one inside and one outside). For such edges, an intersection point is computed to represent the intersection of the reconstructed isosurface with the edge. As the cell graph structure is 3-regular, provided the intersection point lies within the edge’s open interval (ensured numerically by clamping SDF values to non-zero values), the isosurface will intersect the cell into 3-regular polyhedra, with each being split into hexes via midpoint subdivision. In (chernyaev1995marching), thirty-three unique cases are identified based on the possible isosurface topologies within a cube. However, this classification contains a repetition between Case 12.2 and Case 12.3. Furthermore, if one considers mirror symmetry, Cases 11 and 14 are topologically identical. Therefore in total, thirty-one distinct cases remain for analysis.
@@ 99 $(s_{9},0,1)$ $s_{9}(x_{6},y_{6},z_{6})+(1-s_{9})(x_{5},y_{5},z_{5})$
For the remaining seventeen cases (left picture of each case, Figure 7), the concave geometry of the resulting meshes makes such guarantees difficult to establish. While exploring less restrictive constraints on vertices is a potential direction for future work, it is found that aside from Cases 7-3 and 13-4, the other fifteen cases are exceptionally difficult to maintain positive Jacobians within one cell. Consequently, from an implementation standpoint, as shown in the middle picture of each case in Figure 7, a local refinement lookup table is established. For each case, the left picture shows the triangulated isosurface to be reconstructed. The middle picture is the proposed local refinement method, where vertices inside the cell are assigned a virtual inside/outside status (represented by the presence or absence of a black dot). It can be verified that after local refinement, all cells fall into one of the fourteen topologies with positive Jacobian proofs. The right picture displays the reconstructed isosurface, with red dots indicating intersection points on sign-change edges. The reconstructed isosurface preserves the same topology as the triangulated version shown in the left picture. In this way, by refining the cell into multiple cells, the MCHex method achieves greater degree of freedom to represent these concave isosurfaces at the expense of a slight reduction in mesh quality. An important concern arises: in the physical space, to ensure that the locally refined cell shapes can be enumerated and guaranteed to produce positive Jacobians, the newly inserted interior vertices are placed at fixed positions. The true inside/outside status at these points can be different from the inside/outside status. To better understand this issue, in the bottom right picture of Figure 7, the true triangle surface of Case 4-1 is overlaid on the local refinement template. According to the true SDF values, the top-left and bottom-right points should not have black dots, which contradicts the template. This is referred to as sign mismatching. As a result, the reconstructed isosurface may be topologically correct but geometrically inaccurate. The accurate boundary recovery can only be deferred to subsequent heuristic optimization (without guarantee on the upper bounds of mesh quality or geometry accuracy). In the implementation, SDF values for inserted vertices are first computed. When the computed sign contradicts the template requirement, the SDF values are adjusted to a minimal magnitude with the correct sign (e.g., modifying a computed value of $1$ to $10^{-10}$ when the template requires the SDF value to be negative). This strategy maintains maximum fidelity to actual SDF values while only altering signs when strictly necessary. In summary, topological recovery is provided for all seventeen local refinement cases without any ambiguities (zhang2012dual, zhang2012resolving), but completely accurate boundary reconstruction for these cases remains a subject for future work.
@@
Figure 8: Four types of distinct transition cells proposed by (schneiders2000octree). The enumerated cells represent the maximal set in which no two cells are affinely equivalent under orientation-preserving (positive-determinant) transformations.
@@
This work considers the four transition cells introduced in (schneiders2000octree). Figure 8 illustrates these four configurations handling face transitions as well as transitions across convex corners and edges. To reduce redundant validation effort, Theorem 5.2 is introduced to avoid repeatedly verifying the Jacobian for cells that are affine transformations of already validated ones.
@@ 6 Results and Applications
To validate the MCHex approach, a C++ prototype is implemented following Algorithm 1 and evaluated on the Thingi10K dataset (zhou2016thingi10k) through batch processing. The boundary recovery performance is quantified using the IoU and the HR, which measure the volumetric alignment between the input geometry $\text{Int}(T_{\text{geom}})$ and the hex mesh boundary $\text{Int}(\partial H_{\text{Int}(T_{\text{geom}})})$, as well as the surface alignment between $T_{\text{geom}}$ and $\partial H_{\text{Int}(T_{\text{geom}})}$. A comparative analysis between the MCHex and the RO method is conducted to demonstrate the superior boundary convergence rate of MCHex. For each model, an SDF is precomputed on 3-refinement grids. This SDF is then used by both methods to extract the interior hex mesh $H_{\text{Int}(T_{\text{geom}})}$, where in the RO baseline, a cell is included if the sum of the SDF values at its eight vertices is positive.
@@
Figure 11: MeshGems fails to maintain all positive Jacobians on some simple geometries in Thingi10K (zhou2016thingi10k). Six of them are selected and hex-meshing methods are compared using them. In the first two rows, models contain inconsistent surface normals, and inward-facing triangles are highlighted in red. The third and fourth rows contain self-intersecting triangles, also marked in red. The fifth and sixth rows consist entirely of non-manifold edges; thus, all triangles adjacent to such edges are colored red. For each model, hex meshes are generated with comparable vertex counts using MeshGems, (gao2019feature), MCHex, and MCHex with smoothing (tong2025fast), while ensuring that MCHex uses slightly fewer vertices than the former two. Each method is evaluated using the minimum scaled Jacobian, IoU, and HR. Despite being constrained by 3-refinement templates to guarantee all-positive Jacobians and lack of sharp edge preservation due to the intrinsic nature of the MC approach, MCHex achieves better boundary fitting in most cases, with fewer vertices.
@@
MCHex successfully generates all-hex meshes with strictly positive Jacobians on 100% of the geometries tested. To benchmark its robustness against state-of-the-art hex meshing tools, a comparison is made with two significant studies in the past: one utilizing the commercial software MeshGems (schneider2022large), and another is (gao2019feature). In the MeshGems study, 4,050 hex meshes are produced out of 10,000 attempts; among them, 25 meshes contain invalid elements with negative Jacobians. Due to the slow computational speed of (gao2019feature), which could take anywhere from several minutes to days per model, a full-dataset comparison is infeasible. For the 25 models that produced negative Jacobians with MeshGems, hex meshes of comparable scale using (gao2019feature) and MCHex are generated, with particular attention to ensuring that the vertex count of MCHex is slightly lower than that of the corresponding MeshGems mesh. The resulting MCHex meshes are further smoothed using (tong2025fast). Due to the slow convergence of the optimization algorithm in complex meshes, the process is terminated for some meshes before full convergence is reached. A comprehensive comparison is then conducted, evaluating the number of vertices, minimum scaled Jacobian, IoU, and HR metrics.
@@
As illustrated in Figure 11, six representative examples are selected from the 25 models for visualization. In the triangular surface meshes of row 1 (#135757) and row 2 (#338508), outward-facing triangles are colored white, while inward-facing ones are marked red. It is hypothesized that the inconsistent normal directions caused issues for the first two methods, which rely on dihedral angle-based subdivision strategies, resulting in their failure to produce valid hex meshes. In contrast, MCHex uses SDF interpolation error to guide subdivision, making it robust to normal noise and thus capable of generating valid hex meshes. When the vertex budget is low, MCHex exhibits weaker surface approximation capability than (gao2019feature) due to the 3-refinement scheme. However, as the vertex count increases, MCHex surpasses the latter, since (gao2019feature) encounters fitting problems along red-white triangle boundaries. MeshGems performed poorly, largely because it failed to mesh the left part.
@@
In the triangular surface meshes of row 3 (#1508789) and row 4 (#1004827), self-intersecting triangles are colored red. These self-intersections lead to ambiguities in inside/outside classification, which likely caused MeshGems to fail in capturing most of the geometry in row 4. (gao2019feature) operates by identifying and meshing each connected component independently, partitioned the model in row 3 into over 1,000 separate components. This resulted in a hex mesh with an excessively high vertex count and self-intersections. In contrast, MeshGems and MCHex process the input as a single connected volume, thereby avoiding such self-intersections in the resulting hex mesh. However, for MCHex, certain regions where one component deeply penetrates another may be interpreted as interior by the SDF. As a consequence, the distance from these points to the hex mesh surface cannot be effectively reduced through increasing vertex budget. This explains why the IoU metric improves with increasing vertex count in the plots, while the HR remains unchanged. In row 4, (gao2019feature) produces distorted results, likely due to erroneous inside/outside judgments. By leveraging the robustness of the SDF representation, MCHex is able to generate better hex meshes.
@@
In the triangular surface meshes of row 5 (#138102) and row 6 (#138180), all edges are non-manifold, and consequently, every triangle adjacent to these edges is colored red. The non-manifoldness arises from multiple definitions of the same triangular face, leading to inconsistent inside/outside determinations in MeshGems and resulting in poor boundary fitting. In contrast, both (gao2019feature) and MCHex correctly distinguish interior from exterior.
@@
Figure 12: Under a 10-hour time budget, MCHex generates hex meshes on complex models from the Thingi10K dataset with both positive Jacobian and IoU > 0.9, while MeshGems and (gao2019feature) fail. Twenty-two meshes are shown, with input geometries: #866925, #76475, #135807, #44375, #114029, #1041448, #1074637, #940414, #551075, #321050, #136074, #252784, #669972, #147736, #55280, #888464, #117641, #87688, #97866, #498974, #971425, and #372114.
The true strength of MCHex lies in its significant expansion of the geometric domain amenable to hex meshing. Within the Thingi10K dataset, numerous models exist for which neither MeshGems nor (gao2019feature) can generate satisfactory meshes. As shown in Figure 12, these two methods fail to produce hex meshes on exhibited models, within a 10-hour time limit, while the MCHex method succeeds and simultaneously achieves an IoU greater than 0.9 (a reasonable threshold for boundary fitting accuracy) and guarantees all-positive Jacobians.
@@
Table 2: Statistics of the raw, smoothed (using (tong2025fast)), and simplified (using (gao2017robust)) MCHex models in Figure 13. Each cell contains three values separated by “/”, corresponding to the raw, smoothed, and simplified meshes, respectively.
@@
The hex meshes generated by the MCHex method can be further improved through post-processing techniques such as geometric smoothing (tong2025fast) or singularity structure simplification (gao2017robust). For demonstration, another three relatively simple examples that MeshGems fails to maintain positive Jacobians are selected and presented in Figure 13. The corresponding mesh statistics are summarized in Table 2. Here, a regular point is defined as an interior point adjacent to eight hexes or a boundary point adjacent to four hexes. Similarly, a regular edge refers to an interior edge incident to four hexes or a boundary edge incident to two hexes. After smoothing, the minimum scaled Jacobian shows significant improvement, and the edge ratio is effectively reduced. The IoU also increases. The proportion of regular points and edges remains unchanged, as smoothing only warps the vertex positions without altering the mesh topology. In contrast, after singularity simplification, the minimum scaled Jacobian decreases slightly but remains acceptable. The edge ratio is further reduced in the first two models due to the removal of transition cells. In the third model, however, the edge ratio remains similar since the original mesh is already nearly uniform. Meanwhile, the proportion of regular points and edges increases in all cases, indicating a more structured singularity structure. It is worth noting that the simplification process is computationally expensive, typically taking over ten hours to complete. Moreover, when applied to more complex models, the robustness of the simplification procedure becomes an issue. For instance, in the third model, the simplified mesh exhibits zigzag artifacts along originally smooth boundary curves, a problem that becomes more serious in more complex cases and often leads to simplification failures.

### v3

@@ 4 Properties of 3-regular graph
This section begins by establishing the definitions of a $k$-regular graph, a $k$-regular polyhedron, and the midpoint subdivision operation. It then provides a formal proof that the midpoint subdivision of a polyhedron induces all-hex elements if and only if the polyhedron is 3-regular. The necessity of this condition is also proved by showing that any 3-regular polyhedron satisfy the Global Necessary Condition for hex meshing proposed in (26).
@@
As an extension, it holds that the three-dimensional CW complex $P^{\prime}=(V^{\prime},E^{\prime},F^{\prime},C^{\prime})$ induced by 3-regular two-dimensional CW complex $P=(V,E,F)$ satisfy the hex-meshing Global Necessary Condition (26). This equation establishes a global constraint that the resulting hex mesh in $P^{\prime}$ must satisfy:
@@ 5 Positive Jacobian Proof
This section begins by defining the Jacobian determinant as a key measure of element quality. The MC method (9) is then applied to reconstruct the isosurface within each cell $h\in H_{\text{Int}(T_{\text{AABB}})}$ where the SDF values at its eight grids exhibit sign changes. It is demonstrated that the cells intersected by the isosurface are all 3-regular, making them hex meshable via midpoint subdivision. Among a total of thirty-one topological cases in a marching cube, fourteen templates guaranteed to generate positive Jacobian hexes under midpoint subdivision are listed and proven. The other seventeen cases for which a guarantee failed to be established are transferred into the fourteen templates using local refinement. Local refinement is introduced to transform these failed cases into the fourteen cases with proof certificates. Additionally, since the cell intersected by the isosurface can be transition cells or non-cubic cells, the positive Jacobian proofs are therefore given to not only the cubic cell, but all possible deformed hex elements. As a proof of concept, complete proofs for the vanilla 3-refinement templates presented in (39, 48) are provided. Based on the experience, it is highly feasible to extend the proofs to element-saving 3-refinement templates (21) and 2-refinement templates (51, 33) as well. Such tasks are left for future work.
@@
While a bijective mapping requires a positive Jacobian determinant $\lvert\mathbf{J}\rvert>0$ everywhere within an element, it remains an open problem to check the global minimum in the parameter space. An alternative approach, popularized by the visualization tool ParaView (3) and recent literature (6, 15, 17), is to evaluate the Jacobian at nine points: the eight corners and the element center in the parametric space. The minimum value from these nine points is then used as a metric for element quality. The objective of the following analysis is to provide guarantees for positivity at these nine points.
@@
This proof extends the univariate case established in (29) Appendix B to the multivariate setting. Theorem 5.1 tells that the coefficients $\beta_{\mathbf{i}}$ of the Bernstein basis are related to the monomial coefficients $J_{\mathbf{i}}$ via the linear transformation.
@@
Given an all-hex tessellation of the bounding box $T_{\text{AABB}}$ and the SDF value at eight corners of each cell $h\in H_{\text{Int}(T_{\text{AABB}})}$ with respect to the input geometry $T_{\text{geom}}$, the MC algorithm reconstructs an isosurface. An edge is a sign-change edge if its two endpoints lie on opposite sides of the isosurface (one inside and one outside). For such edges, an intersection point is computed to represent the intersection of the reconstructed isosurface with the edge. As the cell graph structure is 3-regular, provided the intersection point lies within the edge’s open interval (ensured numerically by clamping SDF values to non-zero values), the isosurface will intersect the cell into 3-regular polyhedra, with each being split into hexes via midpoint subdivision. In (9), thirty-three unique cases are identified based on the possible isosurface topologies within a cube. However, this classification contains a repetition between Case 12.2 and Case 12.3. Furthermore, if one considers mirror symmetry, Cases 11 and 14 are topologically identical. Therefore in total, thirty-one distinct cases remain for analysis.
@@ 99 $(s_{9},0,1)$ $s_{9}(x_{6},y_{6},z_{6})+(1-s_{9})(x_{5},y_{5},z_{5})$
For the remaining seventeen cases (left picture of each case, Figure 7), the concave geometry of the resulting meshes makes such guarantees difficult to establish. While exploring less restrictive constraints on vertices is a potential direction for future work, it is found that aside from Cases 7-3 and 13-4, the other fifteen cases are exceptionally difficult to maintain positive Jacobians within one cell. Consequently, from an implementation standpoint, as shown in the middle picture of each case in Figure 7, a local refinement lookup table is established. For each case, the left picture shows the triangulated isosurface to be reconstructed. The middle picture is the proposed local refinement method, where vertices inside the cell are assigned a virtual inside/outside status (represented by the presence or absence of a black dot). It can be verified that after local refinement, all cells fall into one of the fourteen topologies with positive Jacobian proofs. The right picture displays the reconstructed isosurface, with red dots indicating intersection points on sign-change edges. The reconstructed isosurface preserves the same topology as the triangulated version shown in the left picture. In this way, by refining the cell into multiple cells, the MCHex method achieves greater degree of freedom to represent these concave isosurfaces at the expense of a slight reduction in mesh quality. An important concern arises: in the physical space, to ensure that the locally refined cell shapes can be enumerated and guaranteed to produce positive Jacobians, the newly inserted interior vertices are placed at fixed positions. The true inside/outside status at these points can be different from the inside/outside status. To better understand this issue, in the bottom right picture of Figure 7, the true triangle surface of Case 4-1 is overlaid on the local refinement template. According to the true SDF values, the top-left and bottom-right points should not have black dots, which contradicts the template. This is referred to as sign mismatching. As a result, the reconstructed isosurface may be topologically correct but geometrically inaccurate. The accurate boundary recovery can only be deferred to subsequent heuristic optimization (without guarantee on the upper bounds of mesh quality or geometry accuracy). In the implementation, SDF values for inserted vertices are first computed. When the computed sign contradicts the template requirement, the SDF values are adjusted to a minimal magnitude with the correct sign (e.g., modifying a computed value of $1$ to $10^{-10}$ when the template requires the SDF value to be negative). This strategy maintains maximum fidelity to actual SDF values while only altering signs when strictly necessary. In summary, topological recovery is provided for all seventeen local refinement cases without any ambiguities (52, 53), but completely accurate boundary reconstruction for these cases remains a subject for future work.
@@
Figure 8: Four types of distinct transition cells proposed by (39). The enumerated cells represent the maximal set in which no two cells are affinely equivalent under orientation-preserving (positive-determinant) transformations.
@@
This work considers the four transition cells introduced in (39). Figure 8 illustrates these four configurations handling face transitions as well as transitions across convex corners and edges. To reduce redundant validation effort, Theorem 5.2 is introduced to avoid repeatedly verifying the Jacobian for cells that are affine transformations of already validated ones.
@@ 6 Results and Applications
To validate the MCHex approach, a C++ prototype is implemented following Algorithm 1 and evaluated on the Thingi10K dataset (55) through batch processing. The boundary recovery performance is quantified using the IoU and the HR, which measure the volumetric alignment between the input geometry $\text{Int}(T_{\text{geom}})$ and the hex mesh boundary $\text{Int}(\partial H_{\text{Int}(T_{\text{geom}})})$, as well as the surface alignment between $T_{\text{geom}}$ and $\partial H_{\text{Int}(T_{\text{geom}})}$. A comparative analysis between the MCHex and the RO method is conducted to demonstrate the superior boundary convergence rate of MCHex. For each model, an SDF is precomputed on 3-refinement grids. This SDF is then used by both methods to extract the interior hex mesh $H_{\text{Int}(T_{\text{geom}})}$, where in the RO baseline, a cell is included if the sum of the SDF values at its eight vertices is positive.
@@
Figure 11: MeshGems fails to maintain all positive Jacobians on some simple geometries in Thingi10K (55). Six of them are selected and hex-meshing methods are compared using them. In the first two rows, models contain inconsistent surface normals, and inward-facing triangles are highlighted in red. The third and fourth rows contain self-intersecting triangles, also marked in red. The fifth and sixth rows consist entirely of non-manifold edges; thus, all triangles adjacent to such edges are colored red. For each model, hex meshes are generated with comparable vertex counts using MeshGems, (15), MCHex, and MCHex with smoothing (43), while ensuring that MCHex uses slightly fewer vertices than the former two. Each method is evaluated using the minimum scaled Jacobian, IoU, and HR. Despite being constrained by 3-refinement templates to guarantee all-positive Jacobians and lack of sharp edge preservation due to the intrinsic nature of the MC approach, MCHex achieves better boundary fitting in most cases, with fewer vertices.
@@
MCHex successfully generates all-hex meshes with strictly positive Jacobians on 100% of the geometries tested. To benchmark its robustness against state-of-the-art hex meshing tools, a comparison is made with two significant studies in the past: one utilizing the commercial software MeshGems (37), and another is (15). In the MeshGems study, 4,050 hex meshes are produced out of 10,000 attempts; among them, 25 meshes contain invalid elements with negative Jacobians. Due to the slow computational speed of (15), which could take anywhere from several minutes to days per model, a full-dataset comparison is infeasible. For the 25 models that produced negative Jacobians with MeshGems, hex meshes of comparable scale using (15) and MCHex are generated, with particular attention to ensuring that the vertex count of MCHex is slightly lower than that of the corresponding MeshGems mesh. The resulting MCHex meshes are further smoothed using (43). Due to the slow convergence of the optimization algorithm in complex meshes, the process is terminated for some meshes before full convergence is reached. A comprehensive comparison is then conducted, evaluating the number of vertices, minimum scaled Jacobian, IoU, and HR metrics.
@@
As illustrated in Figure 11, six representative examples are selected from the 25 models for visualization. In the triangular surface meshes of row 1 (#135757) and row 2 (#338508), outward-facing triangles are colored white, while inward-facing ones are marked red. It is hypothesized that the inconsistent normal directions caused issues for the first two methods, which rely on dihedral angle-based subdivision strategies, resulting in their failure to produce valid hex meshes. In contrast, MCHex uses SDF interpolation error to guide subdivision, making it robust to normal noise and thus capable of generating valid hex meshes. When the vertex budget is low, MCHex exhibits weaker surface approximation capability than (15) due to the 3-refinement scheme. However, as the vertex count increases, MCHex surpasses the latter, since (15) encounters fitting problems along red-white triangle boundaries. MeshGems performed poorly, largely because it failed to mesh the left part.
@@
In the triangular surface meshes of row 3 (#1508789) and row 4 (#1004827), self-intersecting triangles are colored red. These self-intersections lead to ambiguities in inside/outside classification, which likely caused MeshGems to fail in capturing most of the geometry in row 4. (15) operates by identifying and meshing each connected component independently, partitioned the model in row 3 into over 1,000 separate components. This resulted in a hex mesh with an excessively high vertex count and self-intersections. In contrast, MeshGems and MCHex process the input as a single connected volume, thereby avoiding such self-intersections in the resulting hex mesh. However, for MCHex, certain regions where one component deeply penetrates another may be interpreted as interior by the SDF. As a consequence, the distance from these points to the hex mesh surface cannot be effectively reduced through increasing vertex budget. This explains why the IoU metric improves with increasing vertex count in the plots, while the HR remains unchanged. In row 4, (15) produces distorted results, likely due to erroneous inside/outside judgments. By leveraging the robustness of the SDF representation, MCHex is able to generate better hex meshes.
@@
In the triangular surface meshes of row 5 (#138102) and row 6 (#138180), all edges are non-manifold, and consequently, every triangle adjacent to these edges is colored red. The non-manifoldness arises from multiple definitions of the same triangular face, leading to inconsistent inside/outside determinations in MeshGems and resulting in poor boundary fitting. In contrast, both (15) and MCHex correctly distinguish interior from exterior.
@@
Figure 12: Under a 10-hour time budget, MCHex generates hex meshes on complex models from the Thingi10K dataset with both positive Jacobian and IoU > 0.9, while MeshGems and (15) fail. Twenty-two meshes are shown, with input geometries: #866925, #76475, #135807, #44375, #114029, #1041448, #1074637, #940414, #551075, #321050, #136074, #252784, #669972, #147736, #55280, #888464, #117641, #87688, #97866, #498974, #971425, and #372114.
The true strength of MCHex lies in its significant expansion of the geometric domain amenable to hex meshing. Within the Thingi10K dataset, numerous models exist for which neither MeshGems nor (15) can generate satisfactory meshes. As shown in Figure 12, these two methods fail to produce hex meshes on exhibited models, within a 10-hour time limit, while the MCHex method succeeds and simultaneously achieves an IoU greater than 0.9 (a reasonable threshold for boundary fitting accuracy) and guarantees all-positive Jacobians.
@@
Table 2: Statistics of the raw, smoothed (using (43)), and simplified (using (14)) MCHex models in Figure 13. Each cell contains three values separated by “/”, corresponding to the raw, smoothed, and simplified meshes, respectively.
@@
The hex meshes generated by the MCHex method can be further improved through post-processing techniques such as geometric smoothing (43) or singularity structure simplification (14). For demonstration, another three relatively simple examples that MeshGems fails to maintain positive Jacobians are selected and presented in Figure 13. The corresponding mesh statistics are summarized in Table 2. Here, a regular point is defined as an interior point adjacent to eight hexes or a boundary point adjacent to four hexes. Similarly, a regular edge refers to an interior edge incident to four hexes or a boundary edge incident to two hexes. After smoothing, the minimum scaled Jacobian shows significant improvement, and the edge ratio is effectively reduced. The IoU also increases. The proportion of regular points and edges remains unchanged, as smoothing only warps the vertex positions without altering the mesh topology. In contrast, after singularity simplification, the minimum scaled Jacobian decreases slightly but remains acceptable. The edge ratio is further reduced in the first two models due to the removal of transition cells. In the third model, however, the edge ratio remains similar since the original mesh is already nearly uniform. Meanwhile, the proportion of regular points and edges increases in all cases, indicating a more structured singularity structure. It is worth noting that the simplification process is computationally expensive, typically taking over ten hours to complete. Moreover, when applied to more complex models, the robustness of the simplification procedure becomes an issue. For instance, in the third model, the simplified mesh exhibits zigzag artifacts along originally smooth boundary curves, a problem that becomes more serious in more complex cases and often leads to simplification failures.